A force of 50N exerted over a distance of 3.0m causes a block having a mass of 20kg to move 1.0m find the AMA; the IMA; efficiency.


1
Expert's answer
2021-06-18T11:22:11-0400

Find the actual mechanical advantage:


"AMA=\frac{F_1}{F_2}=\frac{50}{20·9.8}=0.26."

The ideal mechanical advantage:


"IMA=\frac{D_1}{D_2}=\frac13=0.33."

Efficiency:


"\eta=\frac{AMA}{IMA}=0.77."